Numerical aperture of fibre is 0.5 and the core refractive index is 1.48 find the refractive index of the cladding acceptance angle
1
Expert's answer
2018-05-13T08:49:54-0400
n_clad = sqrt( (n_core)^2 - NA^2 ) = sqrt(1.48^2 - 0.5^2) = 1.39